OpenStudy (anonymous):

write an equation of the perpendicular bisector of the line segment whose endpoints are (-1,1) and (7,-5)

OpenStudy (anonymous):

can someone help me ?

OpenStudy (anonymous):

Find the slope between those two points. The line perpendicular to that line will be the will have a slope that is the reciprocal of the slope between the original two points.

OpenStudy (anonymous):

i tried it and was having difficulty

OpenStudy (anonymous):

fist find the midpoint of the line it is \[(\frac{-1+7}{2},\frac{1-5}{2})=(3,-2)\]

OpenStudy (anonymous):

okay that is what I got for the mid point

OpenStudy (anonymous):

then find the slope of the line it is \[\frac{1+5}{-1-7}=\frac{6}{-8}=-\frac{3}{4}\]

OpenStudy (anonymous):

should be (2, -2) as midpoint. and the perpendicular slope is also the non negative reciprocal.

OpenStudy (anonymous):

now you want the slope of the perpendicular line. it is the negative reciprocal of your slope, so change the sign and flip it

OpenStudy (anonymous):

get \(\frac{4}{3}\)

OpenStudy (anonymous):

y=(4/3)x-(8/3)

OpenStudy (anonymous):

yes but it asks for perpendicular so that means you get the negative reciprocal which is \[4/3\]

OpenStudy (anonymous):

now use the point slope formula \[y-y_1=m(x-x_1)\] with \(m=\frac{4}{3}\) and either point you like

OpenStudy (anonymous):

yes you are right, the slope is \(\frac{4}{3}\) for the perpendicular line

OpenStudy (anonymous):

okay so now that we found the slope how do we find the equation

OpenStudy (anonymous):

now we can write \[y-1=\frac{4}{3}(x+1)\]and go from there

OpenStudy (anonymous):

some algebra to finish up

OpenStudy (anonymous):

okay let me do that

OpenStudy (anonymous):

I got y = 4/3x - 5.9 is that right

OpenStudy (anonymous):

how did I do with that. I dont think that this is correct

OpenStudy (anonymous):

is anyone still there